# DEFT: Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time

### v1

@@ III-B1 Discrete Elatic Rods (DER)
DEFORM builds upon DER theory to model three distinct deformation modes of DLOs: bending, twisting, and stretching. DER theory models DLOs using two key families of coordinate frames. First, it introduces Bishop Frames, which provide a twist-free reference state along the centerline of the rod. The Bishop Frame only rotates to follow the curve’s geometry, providing a relaxed baseline configuration for the DLO. Second, the theory employs Material Frames, which describe the actual physical deformation of the rod. Material Frames can be expressed relative to the Bishop Frames through a single angle whose rate of change represents the physical twist of the DLO. To describe the twist of the DLO, DER theory assumes that the angle of the Material Frame with respect to the Bishop Frame is one that minimizes the potential energy of the DLO. We briefly summarize DER theory in this section, but a longer introduction can be found here [24].
@@
Figure 2: An illustration of DER coordinate frames.
Equations of Motion: Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ denote the set of all $n-1$ Material Frames at time $t$. Let $\bm{\alpha}$ represent the vector of material properties for each vertex of the DLO (i.e., mass, bending modulus, and twisting modulus) [4, Section 4.2]. Using these definitions, one can compute the potential energy of the DLO arising from bending and twisting which we denote by $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$. DER theory assumes that the DLO reaches an equilibrium state between each simulation time step to obtain the optimal $\bm{\theta}_{t}^{*}$:
@@
III-B2 Differentiable Discrete Elastic Rods (DDER)
Tuning the DER model parameters, $\bm{\alpha}$, to match real-world DLOs is non-trivial. To address this challenge, DEFORM introduces DDER, which enables the automatic tuning of DLO parameters using real-world data. Predicting $\hat{\textbf{X}}_{t+1}$ with DDER through (7) and (8) involves solving an optimization problem as defined in (5). Consequently, minimizing the prediction loss requires solving a bi-level optimization problem at each time step $t$:
$\underset{\bm{\alpha}}{\min}\hskip 14.22636pt$ $\|\textbf{X}_{t+1}-\hat{\textbf{X}}_{t+1}(\bm{\theta}^{*}_{t})\|_{2}\hskip 42.67912pt$ (9)
$\bm{\theta}^{*}_{t}=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ (10)
Note $\hat{\textbf{X}}_{t+1}$ depends on the previously computed $\bm{\theta}^{*}_{t}$ through (6), (7) and (8). To compute the gradient of (10) with respect to $\bm{\alpha}$, DEFORM employs solvers from Theseus [25], which use PyTorch’s automatic differentiation [26]. As we illustrate via experiments in Section V, this can be a time-consuming process.
III-B3 Integration Method with Residual Learning
To compensate for numerical errors that arise during numerical integration(8), DEFORM incorporates a learning-based framework into the integration method:
$\hat{\textbf{X}}_{t+1}=\hat{\textbf{X}}_{t}+\Delta_{t}(\hat{\textbf{V}}_{t+1}+\text{GNN}(\hat{\textbf{X}}_{t},\hat{\textbf{V}}_{t},\bm{\alpha}))$ (11)
This structure builds on residual learning [27], grounding the neural network in physical laws while leveraging data-driven insights.
III-B4 Inextensiblity Enforcement with Momentum Conservation
After numerical integration, the DLO length may deviate from its initial value due to numerical errors, making it difficult to accurately model inextensible DLOs. A common physics-based approach [13, 14, 15] uses mass-spring systems to minimize this deviation, but this can result in stiff simulations and numerical instability. DEFORM addresses this by enforcing an inextensibility constraint while preserving momentum, leading to more accurate and stable simulation.
III-B5 Challenges of Extending DEFORM to BDLOs
Though DEFORM can accurately simulate single-thread DLOs, it is unsuitable for BDLOs for several reasons. First, the definitions of Bishop and Material Frames do not generalize to BDLOs. To see why this is true, recall that at junctions in a BDLO there are several tangent vectors because there are more than two edges. As a result, satisfying (1) and (2) results in an overdetermined system which makes it impossible to find a rotation matrix $\mathbf{R}^{i}$ that satisfies all constraints. Consequently, the Bishop and Material Frames cannot be defined at these junctions. An alternative to extend DEFORM to BDLOs is to simulate each branch individually using DEFORM. However, this requires repeatedly solving (5), which reduces computational efficiency. This is further complicated when trying to enforce inextensibility using the momentum conservation constraint. To address this, DEFT proposes a computational representation that supports parallel processing, enabling more efficient inference.
Second, relying on automatic differentiation to perform parameter tuning becomes computationally prohibitive when dealing with BDLOs. To address this limitation, DEFT utilizes the analytical gradients of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$. Third, the GNN used by DEFORM during residual learning is tailored to single-strand DLOs. To address this limitation, DEFT introduces a specialized GNN Architecture for residual learning for BDLOs that can compensate for numerical integration errors over long horizon predictions.
@@ III-B5 Challenges of Extending DEFORM to BDLOs
Figure 3: Algorithm Overview of DEFT. In the initialization stage (Section IV-A), DEFT begins by separating the BDLO into a parent DLO and one or more children DLOs. Each DLO is discretized into vertices and represented as elastic rods. This setup allows DEFT to capture the geometric and physical properties required for dynamic simulation. To improve computational efficiency, DEFT then predicts the dynamics of each branch in parallel (Section IV-F). During this process, analytical gradients are provided to solve (10), as detailed in Section IV-C, ensuring efficient and stable convergence. Next, to address numerical errors, DEFT employs a GNN designed to learn the BDLO’s residual dynamics (Section IV-D). By modeling discrepancies between simulated and observed behavior, the GNN refines predictions and enhances overall accuracy. After integration, DEFT enforces constraints (Section IV-E) to enforce physical realism. Inextensibility constraints are applied to each branch, while junction-level constraints ensure proper attachment at branch junctions. Additionally, edge orientation constraints enable the propagation of dynamics across these junctions. Throughout the entire pipeline, all components remain fully differentiable, allowing for efficient parameter learning from real-world data.
IV Methodology
Our objective is to model the dynamic behavior of a BDLO under robotic control. Traditional modeling approaches simplify the problem by assuming that the DLO’s endpoints are each held by separate robot end-effectors [8, 28, 16, 10, 29, 30]. Although this assumption reduces complexity, it limits the potential for dexterous manipulation (e.g., inserting the middle of a BDLO into a plug). This work relaxes this assumption and considers scenarios where the BDLO is grasped at an intermediate point, rather than exclusively at its ends.
This section introduces the key components of BDLO modeling and overviews the various components of DEFT, which are illustrated in Figure 3. The key components of DEFT are: 1) analytical gradients to accelerate the optimization procedure used to solve (5); 2) a GNN architecture tailored for residual learning in BDLO simulations; 3) an inextensible constraint formulation that enables the propagation of dynamics at the junction between branches while conserving momentum, which improves the numerical stability of the simulation; and 4) a parallel programming strategy that ensures efficient simulation and gradient computation during training and simulation. Please note that to simplify the presentation, we omit the time subscript $t$ in Sections IV-C and IV-E.
IV-A Branched Deformable Linear Object Model
Figure 4: An illustration of a branched deformable linear object (BDLO) model. Left: An example configuration of BDLO manipulation. Right: Relevant notation and structure at a junction.
To model a BDLO, we decompose its structure into a hierarchical arrangement of DLOs. The parent DLO is defined as the minimum-length path through the BDLO that connects the two end-effectors (i.e., the points where the structure is manipulated or anchored). All other branches diverging from this primary path are modeled as children DLOs that maintain their physical connection to the parent DLO at their attachment points, which we refer to as junction points. The parent and children DLOs are each modeled as DLOs and inherit all previously introduced notation. For conciseness, we describe the junction point using only three edges at each time instance: two from the parent DLO and one from a child DLO. In principle, the methods in DEFT can be extended to handle junctions with more than three edges.
We distinguish between the parent and child DLOs using superscripts $p$ and $c$, respectively. For instance, if a junction is located at the $i$-th vertex, the associated edges are denoted by $\mathbf{e}^{p,i-1}_{t}$, $\mathbf{e}^{p,i}_{t}$, and $\mathbf{e}^{c,1}_{t}$. An illustration of a BDLO can be found in Figure 4. To capture the influence of bending and twisting, we introduce the moment of inertia of the $i$-th edge which is denoted by $\mathbf{I}^{\text{i}}\in\mathbb{R}^{3\times 3}$. It represents how the segment’s mass distribution affects its resistance to rotational deformations. The orientation of $\mathbf{e}^{i}_{t}$ is denoted by an angle axis vector $\mathbf{\Omega}^{i}_{t}\in\mathbb{R}^{3}$.
Let the orientation and position of the two end effectors at time $t$ be denoted by $\mathbf{u}_{t}\in\mathbb{R}^{12}$. We refer to $\mathbf{u}_{t}$ as the input. We assume that the initial state of the BDLO is fully observable, and that it can be accurately estimated using image segmentation [31] in combination with Gaussian Mixture Models [32].
IV-B DEFT Overview
At each time $t$, DEFT simulates the behavior of a BDLO when given predictions of the vertices and their velocities under a specified input as described in Algorithm 1. It begins by first associating the grasped vertices with the robot end effector inputs $\mathbf{u}_{t}$. DEFT then processes the parent and children DLOs in parallel to compute their respective material frames using (3) (4), which are subsequently used to calculate each branch’s potential energy. DEFT then minimizes this potential energy using analytical gradients to determine the optimal parameters $\bm{\theta}_{t}^{*}$ and updates the potential energy accordingly (Line 2). These forces are integrated forward in time using a Semi-Implicit Euler method combined with residual learning to update the vertex velocities and positions (Lines 3 and 4).
However, due to numerical integration errors, branch lengths may change in length. In addition, up to this point, DEFT treats each branch’s internal dynamics in isolation, without accounting for junction interactions. This omission can hinder the propagation of dynamics across junctions. For example, children DLOs may detach from their parent branches, or deformation effects (e.g., bending and twisting) may not propagate through the junction. To address these issues, DEFT enforces additional constraints to ensure each branch remains inextensible, maintain branch connections, and allow deformation effects to propagate across the junction (Line 5). By applying Algorithm 1 over multiple time increments, DEFT can accurately predict the BDLO’s state over extended time horizons. An algorithm overview is illustrated in Figure 3.
1:Associate grasped vertices with $\textbf{u}_{t}$
2:Calculate Material Frames $\Gamma(\textbf{X}_{t},\bm{\theta}_{t})$triangleright (3),(4)
3:$\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$triangleright Section.IV-C
5:$\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t}+\Delta_{t}\hat{\textbf{V}}_{t+1}+\text{GNN}$triangleright Section.IV-D
6:Constraints enforcement on $\hat{\mathbf{X}}_{t+1}$triangleright Algorithm 2
return $\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}}$
IV-C Analytical Gradient Derivation
The goal of this subsection is to present the analytical gradient of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$. To improve the paper’s readability, we summarize the result in the following theorem whose proof can be found in Appendix -A.
Theorem 1.
Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ be the Material Frame at time $t$, and let $\bm{\alpha}$ contain the bending stiffness $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and twisting stiffness $\beta\in\mathbb{R}$. Suppose the total potential energy is the sum of bending and twisting potential energies:
$\begin{split}P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})\;=\;P_{\text{bend}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\mathbf{B}\bigr{)}+\\ P_{\text{twist}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\beta\bigr{)}.\end{split}$ (12)
Then the partial derivative of the bending and twisting potential energies with respect to $\theta^{i}$ are given by
$\begin{split}\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(&\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\\ &\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},\end{split}$ (13)
$\begin{split}\frac{\partial P_{\mathrm{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\beta^{i}\,\bigl{(}\theta^{i}-\theta^{i-1}\bigr{)}\;-\;\\ \beta^{i+1}\,\bigl{(}\theta^{i+1}-\theta^{i}\bigr{)},\end{split}$ (14)
respectively, where the material curvature $\bm{\omega}^{(i,j)}$ is defined as
$\bm{\omega}^{(i,j)}\;=\;\bigl{(}\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{1}^{j},\;\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{2}^{j}\bigr{)}^{T}\quad\text{for}\quad j\in\{i-1,\,i\},$ (15)
$\overline{\bm{\omega}}^{(i,j)}$ denotes the undeformed material curvature, and the curvature vector $\kappa\mathbf{b}^{i}$ in (15) is
$\kappa\,\mathbf{b}^{i}\;=\;\frac{2\,\bigl{(}\mathbf{e}^{i-1}\times\mathbf{e}^{i}\bigr{)}}{\|\mathbf{e}^{i-1}\|_{2}\,\|\mathbf{e}^{i}\|_{2}\;+\;\mathbf{e}^{i-1}\cdot\mathbf{e}^{i}}.$ (16)
IV-D Integration Method with Residual Learning
This subsection constructs the GNN for residual learning in (11). The motivation for employing a GNN lies in its ability to represent a BDLO as a graph, where each vertex is treated as a node and edges capture both intra-branch connections and junctions. This graph-based formulation translates into an adjacency matrix $\mathbf{A}\in\mathbb{R}^{(n^{p}+n^{c})\times(n^{p}+n^{c})}$, which links individual branch nodes and explicitly encodes node information across branch junctions. This adjacency matrix is defined as:
$\mathbf{A}_{i_{1},i_{2}}=\begin{cases}1&\text{if there is an edge between }i_{1}\text{ and }i_{2},\\ 0&\text{otherwise,}\end{cases}$ (17)
An illustration of $\mathbf{A}$ can be found in Figure 5. In practice, DEFT constructs $\mathbf{A}$ for each BDLO whose behavior a user wants to predict. This adjacency matrix is utilized within a Graph Convolutional Network (GCN) [33] to aggregate learned features. A single GCN layer is defined as
$\text{GCN}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})=\mathbf{\tilde{A}}\,\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\,\mathbf{W},$ (18)
where $\mathbf{\tilde{A}}$ is the normalized adjacency matrix
$\mathbf{\tilde{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (19)
@@ IV-D Integration Method with Residual Learning
Note that $\mathbf{\tilde{A}}$ is the normalization of $\mathbf{A}$ using the degree matrix $\mathbf{D}=\mathrm{diag}(d_{1},d_{2},\ldots,d_{n})$ where $d_{i1}=\sum_{i2}A_{i1,i2}$. This normalization ensures that nodes with a high degree do not disproprtionately influence the overall performance. This ensures consistent feature magnitudes across the learned representation [34]. The node feature matrix, $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\in\mathbb{R}^{(n^{p}+n^{c})\times h}$, is constructed by concatenating its arguments together. The trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ is implemented via a multi-layer perceptron (MLP) [35], which projects input features into latent representations.
Figure 5: The adjacency matrix $\mathbf{A}$ for the junction in Figure 4 captures both self-loops (diagonal entries) and inter-node connections (off-diagonal entries). By embedding each node’s local dynamics alongside its coupling to neighboring nodes, $\mathbf{A}$ enables an graph representation of the BDLO’s behavior.
IV-E Enforcing Constraints
After performing numerical integration, DEFT applies constraints to ensure branch inextensibility, ensure that branches are attached at the junctions, and ensure that deformations are propagated throughout the BDLO. To accomplish this, DEFT updates the positions of each vertex using correction terms $\Delta\hat{\mathbf{X}}_{\text{}}$. One could just update the positions of each vertex by just projecting directly onto the constraint set. However, this induces highly nonlinear changes in momentum and can lead to simulation instability and unrealistic behavior.
Motivated by the PDB method [36], DEFT enforces the aforementioned constraints while trying to preserve momentum. DEFT does this by formulating the computation of the correction term as the solution to an optimization problem. As we illustrate through our experimental evaluation in Section V, this leads to more stable simulations, which corresponds to more realistic predictions. Finally, note that DEFT could formulate the optimization problem to compute the correction terms as a single large optimization problem that considers constraints between all vertices in the BDLO simultaneously. However, this large optimization problem can be difficult to solve rapidly. Instead DEFT enforces these constraints sequentially between neighboring nodes. Next, we summarize each of these optimization problems.
IV-E1 Inextensibility Enforcement
@@
To enforce inextensibility of each branch, we define a constraint function between successive vertices. This function penalizes deviations of each segment’s length from its undeformed length. By applying the following theorem, whose proof can be found in [8, Appendix B.2], one can compute $\Delta\hat{\mathbf{x}}^{i},\Delta\hat{\mathbf{x}}^{i+1}$ explicitly:
@@
$\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (21)
s.t. $\hskip 8.5359pt\Delta\hat{\mathbf{x}}^{j}=0$ $\forall j\notin\{i,i+1\},$
$\hskip 8.5359pt\mathbf{M}^{i}\Delta\hat{\mathbf{x}}^{i}+\mathbf{M}^{i+1}\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{0},$
$\hskip 8.5359pt\mathbf{I}^{i}\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0},$
where $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\in\mathbb{R}^{3}$ is defined in Appendix -B and represents the orientation change of the $i$th edge arising from $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ and $\mathbf{0}$ correspond to the zero vector of appropriate size. The only non-zero elements of the minimizer to this optimization problem can be computed explicitly and are equal to:
@@
IV-E2 Enforcing Attachment at the Junction
To ensure the child branch remains attached to the parent branch, we define a constraint function between theirjunction vertices. By applying the following theorem, whose proof can be found in Appendix -C, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:

### v5

@@ Appendix A Supplementary Material for DER Background
To model bending, twisting, and stretching, DER theory models DLOs using two key families of coordinate frames. First, it introduces Bishop Frames, which provide a twist-free reference state along the centerline of the rod. The Bishop Frame only rotates to follow the curve’s geometry, providing a relaxed baseline configuration for the DLO. Second, the theory employs Material Frames, which describe the actual physical deformation of the rod.
Figure 5: An illustration of DER coordinate frames.
Material Frames can be expressed relative to the Bishop Frames through a single angle whose rate of change represents the physical twist of the DLO. To describe the twist of the DLO, DER theory assumes that the angle of the Material Frame with respect to the Bishop Frame is one that minimizes the potential energy of the DLO. We briefly summarize DER theory in this section, but a longer introduction can be found here [37].
@@
Equations of Motion: DER theory assumes that the DLO reaches an equilibrium state between each simulation time step to obtain the optimal $\bm{\theta}_{t}^{*}$:
@@
Appendix B Supplementary Material for the DEFT Algorithm
B.1 DEFT Contribution Visualization
@@
Figure 6: Overview of DEFT contributions. First, DEFT partitions the BDLO into parent and child DLOs, each discretized as elastic rods for accurate dynamic simulation (§4.1). It then predicts each branch’s dynamics in parallel (§4.5)—using analytical gradients to solve the inner loop for fast, stable convergence (§4.5). A GNN learns residual errors between simulated and observed behavior (§4.3), and constraint modules enforce inextensibility, junction attachments, and edge orientations for physical realism (§4.4). Inextensibility constraints are applied to each branch, while junction-level constraints ensure proper attachment at branch junctions. Additionally, edge orientation constraints enable the propagation of dynamics across these junctions. All components remain fully differentiable, enabling efficient parameter learning from real-world data.
B.2 Definition of $\hat{\bm{\Omega}}^{i}_{\Delta,t}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ and $\hat{\bm{\Omega}}^{i}$
In practice, it can be both challenging and impractical to attach hardware onto each segment for tracking its orientation changes. To address this limitation, we approximate the change in orientation, as illustrated in Figure 7. We define $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ as follows:
$\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\frac{\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}}{||\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}||_{2}}\times\frac{(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})-(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})}{||(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})-(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})||_{2}}$ (18)
Intuitively, the cross product of each edge’s tangent vector provides an approximation of the rotation axis and orientation change for the segment $i$. This approximation becomes more accurate when $\Delta_{t}$ is small.
Next, we describe how to compute $\hat{\bm{\Omega}}^{i}$. Note that it is defined recursively using the computation at the previous time step. As a result just within this appendix, we add a subscript $t$ to the symbol. To compute $\hat{\bm{\Omega}}^{i}_{t+1}$, we first convert both $\hat{\bm{\Omega}}^{i}_{t}$ and the correction rotation $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}}_{t},\Delta\hat{\mathbf{X}}_{t})$ from angle-axis to quaternions. We then update $\hat{\bm{\Omega}}^{i}_{t}$ by applying $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}}_{t},\Delta\hat{\mathbf{X}}_{t})$ in quaternion form, and finally convert the result back to angle-axis coordinates to obtain $\hat{\bm{\Omega}}^{i}_{t+1}$. Note that we do not continuously use quaternions because angle–axis is more convenient for representing angular momentum.
Figure 7: An illustration of $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$
B.3 Residual Learning Details
@@
The node feature matrix $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})$ is constructed by concatenating its arguments together. The trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ is implemented using a multi-layer perceptron (MLP) [38], consisting of two linear layers with a ReLU activation in between and a feature dimension of 64 for message passing. This projects input features into latent representations.
B.4 Supplementary Material for Enforcing Constraints
In what follows, we define $C(\hat{\mathbf{z}})$ for each constraint in Theorem 1, present the solutions as individual theorems with proofs, and summarize their enforcement in Algorithm 2.
B.4.1 Inextensibility Enforcement
@@
To enforce inextensibility of each branch, we define a constraint function between successive vertices. This function penalizes deviations of each segment’s length from its undeformed length. By applying the following theorem, one can compute $\Delta\hat{\mathbf{x}}^{i},\Delta\hat{\mathbf{x}}^{i+1}$ explicitly:
@@
$\underset{\Delta\hat{\mathbf{X}}}{\min}\quad$ $\frac{1}{2}\left(C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (20)
s.t. $\Delta\hat{\mathbf{x}}^{j}=0,\quad\forall j\notin\{i,i+1\},$
$\mathbf{M}^{i}\Delta\hat{\mathbf{x}}^{i}+\mathbf{M}^{i+1}\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{0},$
$\mathbf{I}^{i}\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0}.$
where $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\in\mathbb{R}^{3}$ is defined in Appendix B.2 and represents the orientation change of the $i$th edge arising from $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ and $\mathbf{0}$ correspond to the zero vector of appropriate size. The only non-zero elements of the minimizer to this optimization problem can be computed explicitly and are equal to:
@@
Proof. The proof can be found in [14, Appendix B.2].
B.4.2 Enforcing Attachment at the Junction
To ensure the child branch remains attached to the parent branch, we define a constraint function between their junction vertices. By applying the following theorem, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly: